OpenStudy (anonymous):

a block of mass m  5.00 kg is pulled along a horizontal frictionless floor by a cord that exerts a force of magnitude F  12.0 N at an angle u  25.0°. (a) What is the magnitude of the block’s acceleration? (b) The force magnitude F is slowly increased. What is its value just before the block is lifted (completely) off the floor? (c) What is the magnitude of the block’s acceleration just before it is lifted (completely) off the floor?

OpenStudy (anonymous):

Resolve 12 N force into its components. Horizontal force will be 12cos25. so answer to first part is a=F/m = 12cos25/5

OpenStudy (anonymous):

now vertical component of force is 12sin25. in order to life upwards it must overcome the weight of the body Let the force should be increased to F. then F sin25=5x9.8 and hence F- 49/sin 25. now for third part acceleration can be calculated as a=F/m